5. The body, weighing 4kg, rises at a constant speed v = 2 m / s, under the action of a constant force F. Find the force of force F, as well as the work performed by it during 2s.


1
Expert's answer
2021-04-19T17:08:47-0400

Since the speed of the body is constant, the force "F" should be equal and opposite to the weight of the body. Thus:


"F = mg"

where "m = 4kg" is the mass of the body, and "g = 9.8m\/s^2" is the gravitational acceleration.

The power of this force is then:


"P = Fv = mgv"

where "v = 2m\/s" is the speed of the body.

And the work is:


"W = Pt = mgvt"

where "t = 2s" is the time. Thus, obtain:


"P = 4\cdot 9.8\cdot 2 = 78.4W""W = 4\cdot 9.8\cdot 2\cdot 2 \approx 157J"

Answer. Power: 78.4W, Work: 157J.